# From strong to weak correlations in breathing-mode kagome van der Waals materials: Nb3(F,Cl,Br,I)8 as a robust and versatile platform for many-body engineering

### v1

@@ Abstract
By combining ab initio downfolding with cluster dynamical mean-field theory, we study the degree of correlations in the low-temperature structures of the breathing-mode kagome van der Waals materials Nb3(F,Cl,Br,I)8. We show that the Coulomb correlation strength steadily increases from I to Br, Cl, and F, allowing us to identify Nb3I8 as a weakly correlated (obstructed atomic) insulator whose gap is only mildly affected by the local Coulomb interaction. Nb3Br8 and Nb3Cl8 are strongly correlated insulators, whose gaps are significantly influenced by Coulomb-induced vertex corrections. Nb3F8 is a prototypical bulk Mott-insulator whose gap is initially opened by strong correlation effects. Angle-resolved photoemission spectroscopy measurements comparing Nb3Br8 and Nb3I8 allow us to experimentally confirm these findings by revealing spectroscopic footprints of the degree of correlation. Our calculations further predict that the entire material family can be tuned into correlated charge-transfer or Mott-insulating phases upon electron or hole doping. Our magnetic property analysis additionally shows that inter-layer magnetic frustrations in the high-temperature phase drive the lattice phase transition to the low-temperature structures. The accompanying bilayer hybridization through inter-layer dimerization yields magnetic singlet ground states in the Cl, Br, and I compounds. Our findings establish Nb3X8 as a robust, versatile, and tunable class for van der Waals-based Coulomb and Mott engineering and allow us to speculate on the symmetry-breaking effects necessary for the recently observed Josephson diode effect in NbSe2/Nb3Br8/NbSe2 heterostructures.
@@ I Introduction
Correlation effects resulting from electron-electron (Coulomb) interactions are omnipresent in solids. Depending on the relative Coulomb interaction strength, they manifest in various ways, from renormalization of quasiparticle dispersions in the case of weak correlations [1], via induced shake-off/replica bands in the case of intermediate correlations [2, 3], to completely different ground states in the case of strong correlations [4, 5]. Being able to tune the degree of correlations would therefore allow to engineer many-body properties of quantum materials [6].
@@
Here, we show that the transition metal halide family Nb3X8 holds up to this promise. In their low-temperature phases, these so-called breathing-mode kagome vdW materials [18] host only two bands close to the Fermi level, which are cleanly separated from the rest of the electronic structure as a result of simultaneous in-plane trimerization of the transition metal atoms and out-of-plane dimerization of the vdW layers. Using ab initio downfolding to molecular orbitals, we derive material-specific minimal models, including single-particle (hopping) and many-body (Coulomb) interactions for the entire family Nb3X8, with $\text{X}\in\{\text{F},\text{Cl},\text{Br},\text{I}\}$. Studying them using cluster dynamical mean-field theory [19] shows how the balance between kinetic and Coulomb energy changes from Nb3F8* to Nb3I8. In this way, we prove that Nb3X8 realizes several correlated phases ranging from weakly correlated (obstructed atomic) insulators to strongly correlated Mott insulators, which are furthermore tunable by doping as well as by adjusting the number of involved layers as summarized in Table 1.
Table 1: Electronic phases of bulk and monolayer Nb3X8 for different doping levels. $N_{e}$ labels the number of electrons in the low-energy subspace. For bulk (monolayer) $N_{e}=2$ ($N_{e}=1$) labels the undoped case. $\delta$ refers to small doping with $0<\delta\lesssim 0.5$. The Nb3F8* notations refers to its putative structure (see section II.1. )
[…]
$N_{e}$ Bulk (low temperature)
$2$ 2-band Mott insulator strongly correlated insulators weakly correlated insulator
$2+\delta$ correlated metals metal
$3$ charge-transfer insulators Mott insulators
Monolayer
$1$ 1-band Mott insulators
$1+\delta$ correlated metals
@@
Nb3X8 belongs to the class of layered vdW kagome materials, which are intensively studied due to the interplay of their non-trivial topology and intriguing correlation phenomena [18, 20]. Nb3X8 specifically belongs to the sub-class of breathing-mode kagome materials, with alternating in-plane bond lengths, as illustrated in Fig. 1. Among them, Nb3Cl8 and Nb3Br8 have recently gained much attention due to their promises to represent prototypical single-band or few-band Mott insulators [21, 22, 23, 24], possibly accompanied by strongly correlated magnetic ground states [25, 26, 23]. Nb3Br8 has received further attention due to its role in inducing field-free Josephson diode effects in NbSe2/Nb3Br8/NbSe2 heterostructures [27]. Although some scenarios have been suggested [28], it is not yet clear what microscopic behavior yields the required symmetry breakings and whether it is intrinsic to Nb3Br8 or related to the interfaces.
@@
Experimentally, X-ray diffraction shows a lattice phase transitions in Nb3Cl8 and Nb3Br8 from high- to low-temperature structures, which can be understood as a shift in the vdW layer stacking [29, 30]. Raman spectroscopy [31] corroborates this interpretation but also indicates that phonon symmetries are not affected by the transition. Angle-resolved photoemission spectroscopy (ARPES) data shows gapped electronic structures with rather flat, albeit broadened, bands below the Fermi level for bulk Nb3I8 [32], Nb3Cl8 [33] and Nb3Br8 [34, 24] as well as their admixtures [22]. For Nb3Cl8-xBrx, the magnetic susceptibility further shows a Curie-Weiss behavior at elevated temperatures in the high-temperature phases, accompanied by a sudden drop upon entering the low-temperature phase [29, 30, 35, 25, 33, 22]. This has been interpreted as a transition from a paramagnetic to a non-magnetic phase, and the finite susceptibility at very low temperatures is ascribed to magnetic impurities [29, 30].
@@
Theoretically, the modeling of these materials takes place in several steps. A first impression is given by Density Functional Theory (DFT), which has been used to study the monolayer, bilayer, and bulk structures, in both low- and high-temperature crystal structures [36, 37, 32, 21, 23]. Starting with the monolayer, DFT shows a single, rather flat band crossing the Fermi level, which is half-filled and thus hosts exactly one electron per unit cell. At low temperatures and going to bilayer and bulk structures, there is an alternating strong and weak hybridization between adjacent layers, which leads to a dimerization in the out-of-plane direction [24, 23, 22, 21]. In this case, DFT predicts two rather flat bands around the Fermi level, which show a small, but finite hybridization gap. Since flat bands are indicators for low kinetic energies, the role of competing Coulomb interactions needs to be carefully taken into account. Mean-field DFT calculations are, however, known to underestimate the impact of the latter, especially in the case of $d$ orbitals, which dominate the low-energy space in the case of Nb3X8. Therefore, beyond mean-field theories, which take possibly strong Coulomb interactions accurately into account, are necessary. In the context of Nb3X8, so far this has been done only in the case of monolayer Nb3Cl8 [23, 22, 38], generic bilayer models for Nb3Cl8 and Nb3Br8 [21], as well as simplified models of bulk Nb3Cl8 [23] and Nb3Br8 [24] using higher-level many-body theory, such as dynamical mean-field theory, or alternative supercell approaches [39]. All of these calculations hint towards an intriguing interplay between enhanced Coulomb interactions and kinetic energies, which seem to drive variants of Mott insulating states. For bulk structures, this especially requires a careful analysis of the impact of the out-of-plane dimerization, which has so far only been discussed qualitatively for bilayers of Nb3Br8 and Nb3Cl8 [21] as well as for bulk Nb3Br8 [24].
@@ II Results
[…]
@@ II.1 Nb3X8 Crystal Structure
The breathing-mode kagome structure depicted in Fig. 1 (a) yields an in-plane trimerization of Nb atoms (see small triangles in red). In the out-of-plane direction, there exist two stackings, referred to as the high- (HT) and low-temperature (LT) structures, which are depicted in Fig. 1 (b) together with their Brillouin zones in Fig. 1 (c). The LT stacking leads to a dimerization in the out-of-plane direction with alternating strong and weak hybridization between adjacent layers, as indicated by dashed and solid line arrows in Fig. 1 (b) and discussed in more detail below. Nb3Cl8 and Nb3Br8 undergo a structural phase transition that changes the stacking from the HT to the LT structures at 909090\,90K and 382382382\,382K, respectively [30, 29, 40, 35]. While it is hypothesized that Nb3I8 exhibits a similar transition at even higher temperatures, this has yet to be confirmed experimentally.
@@ II.2 Electronic Structure
II.2.1 Origin of Robust Flat Bands
@@
The formation of flat bands in Nb3X8 plays a crucial role in their electronic and magnetic properties. As the underlying lattice structure is of kagome type, it has been suspected that the flat bands originate from the topologically non-trivial flat bands of the well-known generic kagome band structure as described, e.g., by tight-binding models with isotropic hoppings; see for example Refs. 41, 33. However, as the low energetic orbitals of all Nb3X8 compounds are of predominant Nb $d$ character, this simplified picture has to be reconsidered [42]. Therefore, we calculated the electronic structure of LT Nb3Br8 at different levels of the breathing mode perturbation, which is measured by the ratio $d_{2}/d_{1}$, where $d_{1}$ and $d_{2}$ represent the sizes of the small and large triangles, respectively. When the structure is a conventional kagome lattice ($d_{1}=d_{2}$), the octahedral crystal field causes the $d$ orbitals of the Nb ions to split into the lower-energy $t_{2g}$ ($d_{xy}$, $d_{yz}$, and $d_{xz}$) and the higher-energy $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$) states. In the actual breathing-mode ground state ($d_{2}/d_{1}\approx 1.45$), the octahedral crystal field is distorted, and the $t_{2g}$ states are further split into two sublevels labeled $t_{2g}^{1}$ and $t_{2g}^{2}$.
@@
In Fig. 2 we show the corresponding DFT band structures, starting from the unperturbed conventional kagome lattice ($d_{2}/d_{1}=1$) and indicate Nb $t_{2g}^{1}$ orbitals in red, Nb $t_{2g}^{2}$ orbitals in green, and Nb $e_{g}$ orbitals in blue. From this, we see that in the putative conventional kagome structure, Nb3Br8 would be a metal with various Nb $d$ bands crossing the Fermi level. Upon introducing the breathing mode perturbation, we see that the Nb $t_{2g}^{2}$ bands (start to) fully gap, while the Nb $t_{2g}^{1}$ bands stay behind, nearly pinned to the Fermi level. Finally, in the actual full breathing mode structure ($d_{2}/d_{1}\approx 1.45$), we see that the Nb $t_{2g}^{1}$ bands further split into a set of two very flat bands directly around the Fermi level and a set of four unoccupied bands. Next to this, we also find that the occupied Nb $t_{2g}^{2}$ bands split, yielding two sets of very flat bands, while their unoccupied counterparts only mildly change in their dispersion and now overlap in energy with the Nb $e_{g}$ bands.
@@
The flat band formation in Nb3X8 is thus intimately connected to the Nb trimerization within the breathing mode distortion. In the monolayer limit, we already showed that the Nb $t_{1g}$ band at the Fermi level can be best understood as being formed by a molecular orbital centered at the breathing-mode induced Nb trimer [23]. As such, the origin of the two well-separated low-energy flat bands around the Fermi level in Nb3X8 is very similar to the formation of flat bands in the star-of-David phase of TaS2 [13] or twisted bilayer graphene (TBG) [43]: due to new, larger structures (induced by the CDW in TaS2, the moiré potential in TBG, or here the breathing-mode distortion), new effective orbitals are formed, which are spatially rather far apart from each other. This large spatial separation yields a small effective hopping of electrons between these new effective CDW, moiré, or here molecular orbitals, such that their resulting bands are only weakly dispersive. Simultaneously, gaps in the background electronic structures are opened, leaving (nearly) disentangled flat bands behind. All of this is lifted upon destroying the larger structure, i.e. by melting the CDW in TaS2, tuning away from the magic angle in TBG, or suppressing the breathing mode distortion in Nb3X8.
@@
II.2.2 Molecular Orbital Basis
Figure 3: Nb3X8 DFT Bandstructures. Colors indicate the weight of the anti-bonding orbital $|\psi^{(-)}|^{2}=|\langle\psi^{(-)}|\psi\rangle|^{2}$.

### v2

@@ Abstract
By combining ab initio downfolding with cluster dynamical mean-field theory, we study the degree of correlations in monolayer, bilayer and bulk breathing-mode kagome van der Waals materials Nb3(F,Cl,Br,I)8. Our new material-specific many-body model library shows that in low-temperature bulk structures the Coulomb correlation strength steadily increases from I to F, allowing us to identify Nb3I8 as a weakly correlated insulator, Nb3Br8 and Nb3Cl8 as strongly correlated insulators, and Nb3F8 as a prototypical bulk Mott-insulator. Angle-resolved photoemission spectroscopy measurements comparing Nb3Br8 and Nb3I8 allow us to experimentally confirm these findings by revealing spectroscopic footprints of the degree of correlation. Our calculations uncover how the thickness and the stacking affect the degree of correlations and predict that the entire material family can be tuned into correlated charge-transfer or Mott-insulating phases upon doping. Our magnetic property analysis based on our model parameter library additionally confirms that inter-layer magnetic interactions drive the lattice phase transition to the low-temperature structures. The accompanying bilayer hybridization through inter-layer dimerization yields magnetic singlet-like ground states in the Cl, Br, and I compounds. We further prove that all low-temperature compounds are dynamically stable and that electron-phonon coupling to the low-energy subspace is suppressed. Our findings establish Nb3X8 as a robust, versatile, and tunable class for van der Waals-based Coulomb and Mott engineering with a rich phase diagram and allow us to speculate on the symmetry-breaking effects necessary for the recently observed Josephson diode effect in NbSe2/Nb3Br8/NbSe2 heterostructures.
@@ I Introduction
Correlation effects resulting from electron-electron (Coulomb) interactions are omnipresent in solids. Depending on the relative Coulomb interaction strength, they manifest in various ways, starting from renormalizations of quasiparticle dispersions in the case of weak correlations [1], via induced shake-off/replica bands in the case of intermediate correlations [2, 3], or in form of completely different ground states in the case of strong correlations [4, 5]. Being able to tune the degree of correlations would therefore allow to engineer many-body properties of quantum materials [6].
@@
Here, we show that the transition metal halide family Nb3X8 holds up to this promise. These so-called breathing-mode kagome vdW materials [18] host only one band per monolayer close to the Fermi level, which is cleanly separated from the rest of the electronic structure as a result of in-plane trimerization of the transition metal atoms. In multilayer structures this further competes with an out-of-plane dimerization, especially at low-temperatures.
Using ab initio downfolding to molecular orbitals, we derive material-specific minimal models, including single-particle (hopping) and many-body screened Coulomb interaction matrix elements for the entire family Nb3X8, with $\text{X}\in\{\text{F},\text{Cl},\text{Br},\text{I}\}$ ranging from monolayer, bilayer, and to bulk structures. Studying them using cluster dynamical mean-field theory [19] shows how the balance between kinetic and screened Coulomb interaction energies changes from Nb3F8 to Nb3I8 as well as from monolayer to bulk structures. In this way, we prove that Nb3X8 realizes several correlated phases ranging from weakly correlated band insulators to correlated and strongly correlated Mott insulators, which we summarize in the generalized phase diagram spanned by the in-plane versus out-of-plane correlation strengths, as shown in Fig. 1. We further clarify the distinction between the band insulator and correlated insulator through experimental angle-resolved photoemission spectroscopy and by theoretically investigating their response to charge doping. Finally, we present a fully ab initio investigation of the magnetic ground state properties of all low-temperature bulk compounds and specifically on the high- to low-temperature phase transition in Nb3Cl8 and Nb3Br8. This finally allows us to suggest a new mechanism for time-reversal symmetry-breaking in finite stacks of Nb3X8, which is needed to understand the experimentally observed Josephson diode effect in NbSe2/Nb3Br8/NbSe2 heterostructures.
Figure 1: Phase diagram of Nb3X8 for monolayer, bilayer, and bulk structures in the low-temperature phase. The horizontal axis measures the in-plane correlation strength and the vertical axis the out-of-plane correlation strength with $\tau_{\parallel}$, $\tau_{\perp}$ and $U_{0}$ being material and structure specific effective in- and out-of-plane hopping and local Coulomb interaction matrix elements. The phase-space is divided into Mott, correlated, and band insulators.
@@
Nb3X8 belongs to the class of layered vdW kagome materials, which are intensively studied due to the interplay of their non-trivial electronic band topology and intriguing electron correlation phenomena [18, 20]. Nb3X8 specifically belongs to the sub-class of breathing-mode kagome materials, with alternating in-plane bond lengths, as illustrated in Fig. 2. Among them, Nb3Cl8 and Nb3Br8 have recently gained much attention due to their promises to represent prototypical single-band or few-band Mott insulators [21, 22, 23, 24, 25, 26], possibly accompanied by strongly correlated magnetic ground states [21, 27, 28, 24]. Nb3Br8 has received further attention due to its role in inducing field-free Josephson diode effects in NbSe2/Nb3Br8/NbSe2 heterostructures [29]. Although some scenarios have been suggested [30], it is not yet clear what microscopic behavior yields the required symmetry breakings and whether it is intrinsic to Nb3Br8 or related to the interfaces.
@@
Experimentally, X-ray diffraction shows lattice phase transitions in Nb3Cl8 and Nb3Br8 from high- to low-temperature structures, which can be understood as a shift in the vdW layer stacking [31, 32]. Raman spectroscopy [33] corroborates this interpretation but also indicates that phonon symmetries are not affected by the transition. Angle-resolved photoemission spectroscopy (ARPES) data shows gapped electronic structures with rather flat, albeit broadened, bands below the Fermi level for bulk Nb3I8 [34], Nb3Cl8 [35] and Nb3Br8 [36, 25] as well as their admixtures [23]. Spectral gaps in Nb3Cl8 have also been reported from transport measurements [37, 26]. For Nb3Cl8-xBrx, magnetic susceptibilities further show Curie-Weiss behavior at elevated temperatures in the high-temperature phases, accompanied by sudden drops upon entering the low-temperature phases [31, 32, 38, 27, 35, 23]. This has been interpreted as a transition from a paramagnetic to a non-magnetic phase, and the finite susceptibility at very low temperatures is ascribed to magnetic impurities [31, 32].
@@
Theoretically, the modeling of these materials takes place in several steps. A first impression is given by Density Functional Theory (DFT), which has been used to study the monolayer, bilayer, and bulk structures, in both low- and high-temperature crystal structures [39, 40, 34, 22, 24]. Starting with the monolayer, DFT shows a single, rather flat band crossing the Fermi level, which is half-filled and thus hosts exactly one electron per unit cell. At low temperatures and going to bilayer and bulk structures, there is an alternating strong and weak hybridization between adjacent layers, which leads to a dimerization in the out-of-plane direction [25, 24, 23, 22]. In this case, DFT predicts two rather flat bands around the Fermi level, which show a small, but finite hybridization gap. Since flat bands are indicators for low kinetic energies, the role of competing Coulomb interactions needs to be carefully taken into account. Mean-field DFT calculations are, however, known to underestimate the impact of the latter, especially in the case of $d$ orbitals which dominate the low-energy space in the case of Nb3X8. Therefore, beyond mean-field theories, which take possibly strong Coulomb interactions accurately into account, are necessary. So far this has been done only inconsistently for monolayer Nb3Cl8 [21, 23, 41], generic bilayer models approximating Nb3Cl8 and Nb3Br8 [22], as well as for simplified models of bulk Nb3Cl8 [24] and Nb3Br8 [25] using different variants of higher-level many-body theory or alternative supercell approaches [42]. All of these calculations hint towards an intriguing interplay between enhanced Coulomb interactions and kinetic energies, which seem to drive variants of Mott insulating states. For bulk structures, this especially requires a careful analysis of the impact of the out-of-plane dimerization, which has so far only been discussed qualitatively for bilayers of Nb3Br8 and Nb3Cl8 [22] as well as for bulk Nb3Br8 [25]. Most importantly, the role of screening and its impact to the actual material-specific Coulomb interactions along the Nb3X8 family members has not been studied yet.
@@ II Results
To systematically, material-specifically, and quantitatively study these effects, we calculate and investigate all kinetic and Coulomb interaction terms of Nb3X8 with X $\in$ {F,Cl,Br,I} in their monolayer, bilayer, and bulk forms and solve the resulting models using cluster dynamical mean-field theory. Since the high-temperature structure of Nb3I8 has not been observed yet and since correlation effects are most intriguing in the low-temperature phases, we mostly focus on the latter in the following and will only comment on differences in the high-temperature structures of Nb3Cl8 and Nb3Br8 where appropriate. We further put a special focus on the Coulomb interactions and, most importantly, their screening to uncover the exact microscopic origin of various resulting phases. We apply the same formalism to all studied compounds allowing for direct comparisons.
@@ II.1 Nb3X8 Crystal Structure
The breathing-mode kagome structure depicted in Fig. 2 (a) yields an in-plane trimerization of Nb atoms (see small triangles in red). In the out-of-plane direction, there exist two stackings, referred to as the high- (HT) and low-temperature (LT) structures, which are depicted in Fig. 2 (b) together with their Brillouin zones in Fig. 2 (c). The LT stacking leads to a dimerization in the out-of-plane direction with alternating strong and weak hybridization between adjacent layers, as indicated by dashed and solid line arrows in Fig. 2 (b) and discussed in more detail below. Nb3Cl8 and Nb3Br8 undergo a structural phase transition that changes the stacking from the HT to the LT structures at around 9090\,K and 380380\,K, respectively [32, 31, 43, 38]. While it is hypothesized that Nb3I8 exhibits a similar transition at even higher temperatures, this has yet to be confirmed experimentally.
@@ II.2 Electronic Structure
II.2.1 In-Plane Trimerization as Origin of Robust Flat Bands
@@
Figure 4: Nb3X8 DFT Bandstructures. Colors indicate the weight of the bonding ($\psi^{(+)}$) and anti-bonding ($\psi^{(-)}$) orbitals in the Wannier basis. $\psi^{(\pm)}=|\langle\psi^{(\pm)}|\psi\rangle|^{2}$, where $|\psi\rangle$ is the eigenfunction in the Wannier basis.
@@
The formation of flat bands in Nb3X8 plays a crucial role in their electronic and magnetic properties. As the underlying lattice structure is of kagome type, it has been suspected that the flat bands originate from the topologically non-trivial flat bands of the well-known generic kagome band structure as described, e.g., by tight-binding models with isotropic hoppings; see for example Refs. 44, 35. However, as the low energetic orbitals of all Nb3X8 compounds are of predominant Nb $d$ character, this simplified picture has to be reconsidered [45]. Therefore, we calculated the electronic structure of LT Nb3Br8 at different levels of the breathing mode perturbation and thus degree of in-plane Nb trimerization, which is measured by the ratio $d_{2}/d_{1}$, where $d_{1}$ and $d_{2}$ represent the sizes of the small and large triangles, respectively, cf. Fig. 3a. When the structure is a conventional kagome lattice ($d_{1}=d_{2}$), the octahedral crystal field causes the $d$ orbitals of the Nb ions to split into the lower-energy $t_{2g}$ ($d_{xy}$, $d_{yz}$, and $d_{xz}$) and the higher-energy $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$) states. In the actual breathing-mode ground state ($d_{2}/d_{1}\approx 1.47$), the octahedral crystal field is distorted, and the $t_{2g}$ states are further split into two sublevels labeled $t_{2g}^{1}$ and $t_{2g}^{2}$.
@@
In Fig. 3b we show the corresponding DFT band structures, starting from the unperturbed conventional kagome lattice ($d_{2}/d_{1}=1$) and indicate Nb $t_{2g}^{1}$ orbitals in red, Nb $t_{2g}^{2}$ orbitals in green, and Nb $e_{g}$ orbitals in blue. From this, we see that in the putative conventional kagome structure, Nb3Br8 would be a metal with various Nb $d$ bands crossing the Fermi level. Upon trimerization, i.e. upon introducing the breathing mode perturbation, we see that the Nb $t_{2g}^{2}$ bands open up a full gap, while the Nb $t_{2g}^{1}$ bands stay behind, nearly pinned to the Fermi level. Finally, in the actual full breathing mode structure ($d_{2}/d_{1}\approx 1.47$), we see that the Nb $t_{2g}^{1}$ bands further split into a set of two very flat bands located directly around the Fermi level and a set of four unoccupied bands. Next to this, we also find that the occupied Nb $t_{2g}^{2}$ bands split, yielding two sets of very flat bands, while their unoccupied counterparts only mildly change in their dispersion and now overlap in energy with the Nb $e_{g}$ bands.
@@
The flat band formation in Nb3X8 is thus intimately connected to the Nb trimerization within the breathing mode distortion and independent of the LT or HT stacking. In the monolayer limit, we already showed that the Nb $t_{1g}$ band at the Fermi level can be best understood as being formed by a molecular orbital centered at the breathing-mode induced Nb trimer [24]. As such, the origin of the two well-separated low-energy flat bands around the Fermi level in Nb3X8 is conceptually very similar to the formation of flat bands in the star-of-David phase of TaS2 [13] or twisted bilayer graphene [46]: due to new, larger structures (induced by the CDW in TaS2, the moiré potential in twisted bilayer graphene, or here the breathing-mode distortion), new effective orbitals are formed, which are spatially rather far apart from each other. This large spatial separation yields a small effective hopping of electrons between these new effective CDW, moiré, or here molecular orbitals, such that their resulting bands are only weakly dispersive. Simultaneously, gaps in the rest of the electronic structures are opened, leaving (nearly) isolated flat bands behind. All of this is lifted upon destroying the larger structure, i.e. by melting the CDW in TaS2, tuning away from the magic angle in twisted bilayer graphene, or suppressing the breathing mode distortion in Nb3X8.
@@
II.2.2 Out-of-Plane Dimerization as Origin of Flat Band Gapping